Three times one number plus a second number is 40,and the one number minus the second number is 4. Find the numbers.
Answer by macston(5194)

F=first number: S=second number
F-S=4
F=S+4
3F+S=40 Substitute for F
3(S+4)+S=40
3S+12+S=40
4S=28
S=7 ANSWER 1: The second number is 7.
F=S+4=7+4=11 ANSWER 2: The first number is 11
CHECK:
3F+S=40
3(11)+7=40
33+7=40
40=40